# Advection of the image point in probabilistically-reconstructed phase spaces

### v2

@@ Abstract
Insufficient reference data is ubiquitous in data-driven computational fluid dynamics, as it is usually too expensive to compute or impossible to observe over long enough times needed for data-driven methods to properly operate out of the sampled range. Ultimately, the lack of data can significantly compromise the fidelity of results computed with data-driven methods or make the data-driven approach inapplicable. In order to challenge this problem, we propose a probabilistic reconstruction method which enriches the hyper-parameterisation approach with ideas underlying the probabilistic-evolutionary approach. We offer to use Advection of the image point (a hyper-parameterisation method) on data sampled from the joint probability distribution of the reference dataset. The idea is to blend together the simplicity of the hyper-parameterisation (HP) method with an extra source of reference data provided through sampling from the joint probability distribution.
The HP method has been tested on the sea surface temperature and surface relative vorticity computed with the global 1/4 and 1/12 resolution NEMO model. Our results show that the HP solution (the solution computed with the HP method) in the probabilistically-reconstructed and reduced (in terms of dimensionality) phase space at 1/4 resolution is more accurate compared to the 1/4-solution computed with NEMO and it is also several orders of magnitude faster to compute than the 1/4-run. The proposed method shows encouraging results for the NEMO model and the potential for the use in other operational ocean and ocean-atmospheric models for both deterministic and probabilistic predictions. The method can also be used as a fast reanalysis in which the complex dynamics of a comprehensive ocean model is replaced with the HP solution, as well as a dynamic interpolation method to stitch gaps in observational data.
keywords:
hyper-parameterization , NEMO ocean model , data-driven method , probabilistic reconstruction , phase space , state space , reanalysis , dynamic interpolation
@@ 1 Introduction
The data-driven paradigm is gradually gaining momentum in computational fluid dynamics by offering an alternative to the physics-driven approach in situations when there is no physics-based model for the studied phenomenon or it is too computationally-expensive to acquire data from the model (e.g. [1] and references therein). The data-driven approach includes a variety of methods, spreading across different fields (e.g. statistical modelling [2, 3], dynamical system reconstruction [4, 5, 6, 7], hybrid modelling [8, 9], probabilistic-evolutionary approach [10], etc.) In recent years, machine learning (ML) methods started to enrich the data-driven approach with artificial neural networks (e.g. [11, 12, 13, 14, 15, 16]) typically trained on reanalysis data or reference runs of physics-driven models. The majority of these ML methods are used for predictive modelling (e.g. [17, 18, 19]), while some, known as Interpretable ML, go beyond traditional ML predictions and try to improve our understanding of modelled phenomena (e.g. [20, 21]).
@@
Being a powerful tool providing a wide spectrum of methods ranging from purely-data to physics-data driven, also know as hybrids, the data-driven approach typically suffers from the lack of reference data that can significantly compromise its fidelity or make it inapplicable. In order to help this problem, we offer a probabilistic reconstruction method which augments the hyper-parameterisation (HP) approach (e.g., [22, 8, 23, 24, 9]) with some ideas underlying the probabilistic-evolutionary approach [10]. More specifically, we propose to use Advection of the image point (a purely data-driven HP method working in phase space, also called state space) together with data sampled from the joint probability distribution of reference data so as to probabilistically reconstruct the reference phase space for the HP method and make it operable for short reference data records otherwise falling out of its range of applicability. The main difference between the proposed methodology and ML methods (based on artificial neural networks) is that: the former does not have the training phase (the calibration procedure of the HP method might be interpreted as such though), it uses phase spaces instead of the physical space, respects the regional stability property, and relies on the unconventional calculation of directional vectors. The regional stability is defined as a property of the solution of a differential equation to permanently remain in a neighborhood of the phase space region if the initial condition is sufficiently close to that region. By the unconventional calculation of directional vectors we mean how the right hand side (RHS) in the HP method is calculated. Namely, we do not use a sequence of following each other states from the past (as conventional integrators) to calculate the RHS, but use the states neighboring the current state in the phase space.
The main advantages of the HP method in probabilistically-reconstructed phase spaces are the following:
$\bullet$ Flexibility. The method enables research situations involving HP of only some part/component of the ocean circulation: geographical part (ocean surface or deep layers, different basins, etc.), specific field (salinity, temperature, velocity, etc.), specific observations (drifter, satellites, etc.), or their combination. For example, one can use HP only for the ocean circulation and solve separately for the biochemical processes embedded into it (e.g. carbon cycle research);
$\bullet$ Scalability. HP offers a great resolution scalability and can be used to parametrise the effect of any high-resolution process into lower-resolution systems (e.g. mesoscales processes in larger-scale processes, sub-mesoscale processes in mesoscale-resolving/sub-mesoscale-permitting systems, etc.), and therefore it can be readily adopted by forecasting systems;
$\bullet$ Alternative to parameterisations, which can be used in comprehensive ocean models (NEMO, MITgcm, FESOM, etc.);
$\bullet$ Fast generation of ensembles. HP is well-suited for generating ensembles of solutions, which serve many purposes in ocean and atmospheric modelling;
$\bullet$ HP as a tool for process studies. HP allows one to study oceanic feedbacks on climate. In this case the ocean is fully under control; one can systematically change parameters and properties of the ocean circulation and find out their global consequences. Using HP one can uncover and understand internal links and feedbacks within the system, by replacing specific components with HP analogues.
@@
Although, we consider the use of probabilistically-reconstructed phase spaces within the HP framework, its application is not bound to HP. For example, probabilistically-reconstructed reference datasets can be used within the context of other data-driven methods, as well as for training artificial neural networks, and for reconstruction of missing parts in observational data from different sources (drifters, weather stations, satellites, etc.)
@@ 2 The probabilistic reconstruction method: schematic and basics
The idea underlying the probabilistic reconstruction method is to calculate the joint probability distribution (JPD) for a given reference dataset and then sample reference data from this JPD instead of using the reference dataset itself. Hence, the JPD becomes an extra layer between the reference data and a data-driven method thus providing an extra source of reference data from the reference data distribution. It would be instructive to first look into the schematic of the proposed method (Fig. 1) for better understanding of its specifics and then consider the method in more detail.
@@
(a) Reference data acquisition. The first step is to provide the method with reference data. Depending on the problem to solve, it can be numerical simulations, observations, or both. In our case it is sea surface temperature (SST) computed with a global 1/12 resolution ocean model NEMO and then interpolated onto a 1/4 grid (Fig. 1a); SST is a scalar field defined on a cartesian grid. If observations are available then they can also be included in the reference dataset by simply adding them to the reference data record. In order to be consistent in space, observations must be on the same grid as the NEMO data; consistency in time is not required, as the proposed method does not use it.
(b) Reference phase space calculation. Before proceeding with the second step it would be helpful to remind the reader that a phase space (or state space) is an abstract mathematical space of all possible states of a physical system, while the physical space is typically defined as the actual three-dimensional space we all live in.
@@
The second step is to get the reference phase space for the HP method called ”Advection of the image point” [22, 24]; the reference phase space is shown as an orange blob in Fig. 1b. The latter needs states (which is SST itself) and directional vectors (time derivatives of SST, also referred to as tendencies), which are calculated from the reference data; different schemes can be used, we use the central finite difference in time. The tendencies are shown as black vectors in Fig. 1b, and the points they are attached to represent SST, which is a vector in the phase space, while it is a scalar field in the physical space.
@@
The white spots (which we call voids) are regions of the phase space which lack data (usually, due to short reference records or unreliable observations). Typically, these voids is a reason for why data-driven methods cease to work on meager phase spaces.
@@
(c) Joint Probability Distribution calculation. The third step is to calculate the joint probability distribution (JPD) from the reference data in phase space; it is shown as a blue surface in Fig. 1c for illustration purposes. In high-dimensional phase spaces (like the one used for SST) the JPD is a hyper-surface. It is important to remark that we calculate the JPD only for SST, while SST tendencies are computed in a different way explained below. The JPD for the reference SST is calculated globally, i.e. the whole reference dataset is used in the calculation.
@@
The standard approach to calculate an approximation to the JPD is to use the histogram method. This method is based on the subdivision of the domain into hypercubes (bins) and count the number of states in every bin to compute the height of each column of the multidimensional histogram; the domain is a hypercube in the phase space of SST in our case, the phase space dimension is equal to the number of grid points needed to approximate SST in the physical space. This method works well only in low-dimensional spaces, since the computer memory needed grows as the number of bins to the power of the space dimensions. Therefore, in high-dimensional spaces (like the one used for SST) computing a multidimensional histogram and keeping it in memory for further sampling is an unaffordable option.
@@
In this study, we do not calculate the whole multi-dimensional histogram. Instead, we compute a coordinate-wise probability density function (PDF) for every component of the state (i.e. for every grid point of SST) and then use these PDFs to sample from the JPD; it gives us access to all necessary information for sampling at virtually no cost. Notice that these PDFs can be interpolated if there is not enough reference data to accurately approximate the PDFs.
@@
(d) Sampling from JPD and calculation of probabilistically-reconstructed phase space. In the fourth step we sample SST from the JPD and then calculate SST tendencies by averaging reference SST tendencies over the neighborhood of sampled SST (Fig. 2).
@@
The localized calculation of tendencies allows to decrease the dimensionality of the JPD two times (as the JPD dimension for both SST and its tendencies is two times larger compared to the JPD only for SST) and thus make its calculation more efficient without compromising on accuracy. It is worth mentioning that tendencies can also be computed in a probabilistic way from a local JPD which is constructed based on the information from the neighborhood of the new tendency (red contour in Fig. 2), but this method is slower than the one used in this study.
@@
In order to sample from the JPD we use the inverse transform sampling method [25]. We have also tried the rejection sampling but did not gain that much of a difference in its favour; a sampling method based on a reverse-time stochastic differential equation [26] might be another option instead of the inverse transform sampling. Sampling from the JPD allows us to probabilistically reconstruct the originally sparsely-populated reference phase space by filling it with new states and directional vectors (tendencies) needed to advect the image point (Fig. 1d). Note that the probabilistically-reconstructed reference data can be used not only in the HP method, but in other data-driven methods which suffer lack of data.
@@
We would like to draw the reader’s attention to the fact that the new states generated by sampling from the JPD does not mean new reference states (i.e. the states generated by the underlying reference model, it is the NEMO model in our case), since sampling from the JPD cannot guarantee that sampled states are reference states, and the HP dynamics in probabilistically-reconstructed phase spaces cannot guaranteed it either.
For a more detailed discussion on how the sampling procedure calculates a new state, we consider a two-dimensional case (i.e. a set of points in an XY-plane) for simplicity. In this case we have two coordinate-wise PDFs (one for each dimension). Sampling starts from computing the PDF for the first dimension and its corresponding CDF. Next, we draw a random number $r_{1}\in U[0,1]$, with $U[0,1]$ being a uniform distribution in the unit interval, and find $x_{1}(r_{1})$ with the inverse transform sampling method, where $x_{1}(r_{1})$ is the first coordinate of the new state. In order to calculate the second coordinate of the new state, we compute the PDF for the second dimension given $x_{1}(r_{1})$ and its CDF. Then, we draw another random $r_{2}\in U[0,1]$ and compute $x_{2}(r_{2})$, which is the second coordinate of the new state. In this way we can sample as many states as we need from the reference JPD.
@@
(e) Advection of the image point in probabilistically-reconstructed phase space. In the fifth step, the probabilistically-reconstructed phase space is used by Advection of the image point to calculate the trajectory (red line in Fig. 1e). For the reader’s convenience, we briefly describe how Advection of the image point works. This method falls into the category of data-driven methods from the hyper-parametersation (HP) class which takes advantage of working in phase spaces, as opposed to the conventional methods operating in the physical space. The method has been tested on both idealized and comprehensive ocean models (two-layer quasi-geostrophic model in a channel and MITgcm in the North Atlantic configuration) and showed significant improvements of the HP solution toward the reference one [22, 24]. The HP approach currently umbrellas five different methods (ranging from purely data-driven to hybrids, which combine the data- and physics-driven paradigms). Another striking feature of the HP approach is that its measure of goodness is how close the HP solution is to the reference phase space. To put it another way, the HP approach matches phase spaces (the reference phase space and the phase space where the HP solution evolves), whereas conventional methods match individual trajectories. This measure allows the HP solution to evolve in the neighborhood of the reference phase space, and therefore reproduce the flow dynamics which is very similar to the reference one.
@@
$\frac{d\mathbf{y}}{dt}=\frac{1}{M}\sum\limits_{j{\color[rgb]{0,0,0}\in}\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})+\eta\left(\frac{1}{N}\sum\limits_{i{\color[rgb]{0,0,0}\in}\mathcal{U}_{I}}\mathbf{x}_{i}-\mathbf{y}(t)\right),\quad\mathbf{y}(t_{0})=\mathbf{y}_{0},$ (1)
@@
In a nutshell, the actual dynamics is given by the observed reference tendency $\frac{1}{M}\sum\nolimits_{j{\color[rgb]{0,0,0}\in}\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})$ the imperfect reconstruction of which from the reference data is compensated by nudging towards the observed reference state $\frac{1}{N}\sum\nolimits_{i{\color[rgb]{0,0,0}\in}\mathcal{U}_{I}}\mathbf{x}_{i}$ in the probabilistically-reconstructed phase space. Note that the neighbourhood is computed as the average over $M$ (and $N$ for the nudging term) nearest, in $l_{2}$ norm, to the solution $\mathbf{y}(t)$ points. The choice of the norm and the way the neighborhood $\mathcal{U}(\mathbf{y}(t))$ is calculated is not limited to those used in this study, and can vary depending on what is needed. Our choice is, probably, the simplest, but it serves well the purpose of this work. We dub the set of parameters $\{M,N,\eta\}$ hyper-parameters. The hyper-parameters can be set based on the chosen metric and available data, we will return to their choice later. The interested reader is referred to [22, 24] for a more detailed discussion of the method.
Figure 3: Shown is a sketch of the probabilistically-reconstructed reference phase space (orange blob), vectors $\mathbf{F}(\mathbf{x}_{i})$ (black arrows) pointing out from points $\mathbf{x}_{i}$. For the HP method all these points is a cloud of points which are not ordered in time. The HP solution $\mathbf{y}(t)$ is shown as a red trajectory, and the neighborhood $\mathcal{U}(\mathbf{y}(t))$ is the black contour surrounding the red dot, which is the HP solution $\mathbf{y}(t)$ at time $t$. The red vectors (and also every point they are attached to) within the neighborhood are the $N$ reference fields $\mathbf{x}_{i},\,i\in\mathcal{U}_{I}$ (and $M$ reference tendencies $\mathbf{F}(\mathbf{x}_{i})$) nearest, in $l_{2}$-norm, to the HP solution $\mathbf{y}(t)$. We assume that $\mathcal{U}_{I}=\mathcal{U}_{J}$ and $M=N$, but the method is not bound to this particular choice.
The characteristic feature of the HP method is that it looks at the reference dataset as being a cloud of points which are not ordered in time. This is why we write $\mathbf{x}_{i}$ and $\mathbf{F}(\mathbf{x}_{j})$ instead of $\mathbf{x}(t)$ and $\mathbf{F}(\mathbf{x}(t))$ in equation (1). At every time step, the HP method takes $N$ nearest points $\mathbf{x}_{i}$ (and $M$ tendencies $\mathbf{F}(\mathbf{x}_{j})$) to the solution $\mathbf{y}(t)$ points (Fig. 3). Thus, the HP method never runs out of reference data, as a set of nearest points to the HP solution $\mathbf{y}(t)$ can always be found. The main advantage of using the HP method in probabilistically-reconstructed phase spaces is that the data for the method is sampled from the same distribution as the reference data itself.

### v3

@@ Abstract
Insufficient reference data is ubiquitous in data-driven computational fluid dynamics, as it is usually too expensive to compute or impossible to observe over long enough times needed for data-driven methods. The lack of data can significantly compromise the fidelity of results computed with data-driven methods or render them inapplicable. To challenge this problem, we propose a probabilistic reconstruction method that enhances the hyper-parameterisation (HP) approach with ideas underlying the probabilistic-evolutionary approach. We offer to use the HP method “Advection of the image point” on data sampled from the joint probability distribution of the reference dataset.
The HP method has been tested regionally on the sea surface temperature and surface relative vorticity computed with the global 1/4 and 1/12 resolution NEMO model. Our results show that the HP solution (the solution computed with the HP method) in the probabilistically-reconstructed and reduced (in terms of dimensionality) phase space at 1/4 resolution is more accurate than the 1/4-solution computed with NEMO. Additionally, the HP solution is several orders of magnitude faster to compute than the 1/4 NEMO solution. The proposed method shows encouraging results for the NEMO model and the potential for the use in other operational ocean and ocean-atmospheric models for both deterministic and probabilistic predictions. Furthermore, the method can be used as a fast reanalysis tool allowing the complex dynamics of a comprehensive ocean model to be approximated by the HP solution. It can also function as a dynamic interpolation method to fill gaps in observational data.
@@
Plain Language Summary
The shortage and damage of reference data used for training or calibrating data-driven methods is a common issue in ocean modelling. This problem can greatly undermine the accuracy of the methods or make them unusable. We propose a solution to mitigate this problem and a hyper-parameterisation (HP) method that can utilize short and damaged data from comprehensive ocean models otherwise falling outside its range of applicability. The methodology proposed in this study not only enables HP to work with sparse and damaged data sets but also serves as a fast reanalysis tool; it approximates complex ocean flows with HP solutions, which are several orders of magnitude faster to compute. Additionally, the proposed methodology can be used as a dynamic interpolation technique to fill gaps in observational data for further use in reanalysis and data assimilation.
@@ 1 Introduction
The data-driven paradigm is gradually gaining momentum in computational fluid dynamics by offering an alternative to the physics-driven approach in situations when there is no physics-based model for the studied phenomenon or it is too computationally expensive to acquire data from the model (e.g. [15] and references therein). The data-driven approach includes a variety of methods, spreading across different fields (e.g. statistical modelling [27, 29], dynamical system reconstruction [2, 7, 5, 14], hybrid modelling [20, 24], probabilistic-evolutionary approach [23], etc.) In recent years, machine learning (ML) methods started to enrich the data-driven approach with artificial neural networks (e.g. [18, 31, 16, 30, 11, 6]) typically trained on reanalysis data or reference runs of physics-driven models. The majority of these ML methods are used for predictive modelling (e.g. [12, 4, 13]), while some, known as Interpretable ML, go beyond traditional ML predictions and try to improve our understanding of modelled phenomena (e.g. [1, 3]).
@@
Being a powerful tool providing a wide spectrum of methods ranging from purely-data to physics-data driven, also known as hybrids, the data-driven approach typically suffers from the lack of reference data that can significantly compromise its fidelity or make it inapplicable. In order to help this problem, we offer a probabilistic reconstruction method which augments the hyper-parameterisation (HP) approach (e.g., [19, 20, 21, 22, 24]) with some ideas underlying the probabilistic-evolutionary approach [23]. More specifically, we propose to use Advection of the image point (a data-driven HP method working in phase space, also called state space) together with data sampled from the joint probability distribution of reference data so as to probabilistically reconstruct the reference phase space for the HP method and make it operable for short or damaged (unreliable for further use) reference data records otherwise falling out of its range of applicability. The main advancements of this study relative to the previous HP works are: the introduction of a probabilistic phase-space reconstruction layer based on JPD sampling; a dimensionally efficient strategy that samples only state distributions while computing tendencies locally; demonstration of robustness to severely damaged reference datasets; application in a high-dimensional ocean modelling context via reduced EOF representation. The main difference between the proposed methodology and ML methods (based on artificial neural networks) is that: the former does not have the training phase (the calibration procedure of the HP method might be interpreted as such though), it uses phase spaces instead of the physical space, respects the regional stability property, and relies on the unconventional calculation of directional vectors. The regional stability is defined as a property of the solution of a differential equation to permanently remain in a neighborhood of the phase space region if the initial condition is sufficiently close to that region. By the unconventional calculation of directional vectors we mean how the right-hand side (RHS) in the HP method is calculated. Namely, we do not use a sequence of following every other states from the past (as conventional integrators) to calculate the RHS, but use the states neighboring the current state in the phase space.
@@
Although we consider the use of probabilistically-reconstructed phase spaces within the HP framework, its application is not bound to HP alone. For example, probabilistically-reconstructed reference datasets can be used within the context of other data-driven methods, including the training of artificial neural networks and reconstructing missing parts in observational data from different sources (drifters, weather stations, satellites, etc.)
The term probabilistic, used throughout this study, refers to two distinct but complementary aspects of the method. First, the HP framework is probabilistic in the sense that it is grounded in the empirical estimation of the joint probability distribution (JPD). Local averages used in the method, and optionally second moments, are computed from this distribution using neighborhood-based statistics. While the resulting dynamics is deterministic once the reference phase space is fixed, its structure is governed entirely by the statistical geometry of the data. The evolution is deterministic, but it is driven by empirical estimates derived from a probability distribution, not from first-principles physics. Second, the method is probabilistic in the stricter sense that, when the reference phase space is sparse or contains voids, it can be reconstructed or augmented by sampling from the JPD. This probabilistic filling is performed offline, prior to trajectory integration, and is crucial in enabling the HP method to operate in regions poorly represented by the raw reference data. However, once the reference phase space is built the HP trajectory evolves deterministically. In this way, the probabilistic nature of the method lies in both the statistical geometry of the phase space and the probabilistic reconstruction of under-sampled regions, but not in the runtime dynamics itself.
it ’s also important to note that the method does not explicitly identify voids in phase space. Instead, it relies on sampling from the estimated JPD, which increases density in statistically under-sampled regions. Void filling is therefore implicit and governed by the estimated probability density. Regions with zero empirical support cannot be reconstructed.
@@ 2 The probabilistic reconstruction method: schematic and basics
The idea underlying the probabilistic reconstruction method is to calculate the joint probability distribution (JPD) for a given reference dataset and then sample reference data from this JPD instead of using the reference dataset itself. Hence, the JPD becomes an extra layer between the reference data and a data-driven method thus providing an extra source of reference data from the reference data distribution. It would be instructive to first look into the proposed method’s schematic (figure 1) to better understand its specifics and then consider the method in more detail.
@@
(a) Reference data acquisition. The first step is to provide the method with reference data. Depending on the problem to solve, it can be numerical simulations, observations, or both. In our case it is sea surface temperature (SST) computed with a global 1/12 resolution ocean model NEMO and then interpolated onto a 1/4 grid (figure 1a); SST is a scalar field defined on a cartesian grid. If observations are available then they can also be included in the reference dataset by simply adding them to the reference data record. In order to be consistent in space, observations must be on the same grid as the NEMO data. It is important to note that two distinct roles of time must be clearly separated in the methodology:
1. Preprocessing stage (derivative estimation). The reference tendencies $F(x_{i})$ are estimated using a central finite-difference approximation, which requires temporally ordered data. Temporal consistency is therefore essential during this preprocessing step.
2. HP trajectory reconstruction stage. Once state-tendency pairs $(x_{i},F(x_{i}))$ are constructed, the HP method operates purely in phase space. The reference data are treated as an unordered cloud of points, and chronological ordering is no longer used during trajectory integration.
(b) Reference phase space calculation. Before proceeding with the second step it would be helpful to remind the reader that a phase space (or state space) is an abstract mathematical space of all possible states of a physical system, while the physical space is typically defined as the actual three-dimensional space we all live in. In this study, a discretized scalar field defined on a spatial grid with $n$ grid points is represented as a vector in $\mathbb{R}^{n}$, so that each grid point corresponds to one coordinate of the phase space. This grid-point representation is not essential to the method. More generally, the HP framework operates in an abstract state space and allows for arbitrary coordinate transformations of the data. For example, one may represent the system via EOF-PC decomposition, wavelet transforms, spectral expansions, or other linear or nonlinear mappings. The grid-point formulation used here is therefore a special case of a broader class of admissible phase space representations.
@@
The second step is to get the reference phase space for the HP method called ”Advection of the image point” [19, 22]; the reference phase space is shown as an orange blob in figure 1b. The latter needs states (which is SST itself) and directional vectors (time derivatives of SST, also referred to as tendencies), which are calculated from the reference data; different schemes can be used, we use the central finite difference in time. The tendencies are shown as black vectors in figure 1b, and the points they are attached to represent SST, which is a vector in the phase space, while it is a scalar field in the physical space.
@@
The white spots (which we call voids) are regions of the phase space which lack data (usually, due to short reference records or damaged observations). Typically, these voids are a reason why data-driven methods cease to work on meager phase spaces.
@@
(c) Joint Probability Distribution calculation. The third step is to calculate the joint probability distribution (JPD) from the reference data in phase space; it is shown as a blue surface in figure 1c for illustration purposes. In high-dimensional phase spaces, such as the one used for SST, the JPD is a hyper-surface. It is important to remark that we calculate the JPD solely for SST, while SST tendencies are computed in a different way explained below. The JPD for the reference SST is calculated globally, meaning that the entire reference dataset is used in this calculation.
@@
The standard approach to calculate an approximation to the JPD is to use the histogram method. This method is based on the subdivision of the phase space into hypercubes (bins) and counting the number of states in each bin to compute the height of each column of the multidimensional histogram. The method works well only in low-dimensional spaces, since the computer memory needed grows as the number of bins to the power of the space dimensions. Therefore, in high-dimensional spaces (like the one used for SST) computing a multidimensional histogram and keeping it in memory for further sampling is an unaffordable option.
@@
In this study, we do not calculate the whole multi-dimensional histogram. Instead, we compute a marginal probability density function (PDF), or coordinate-wise PDF, for each component of the state (i.e. for each grid point of SST) and then use these PDFs to sample from the JPD; it gives us access to all necessary information for sampling at virtually no cost. Notice that these PDFs can be interpolated if there is insufficient reference data to accurately approximate them.
@@
(d) Sampling from JPD and calculation of probabilistically-reconstructed phase space. In the fourth step, we sample SST from the JPD and then calculate SST tendencies by averaging reference SST tendencies over the neighborhood of sampled SST (figure 2).
@@
This neighborhood-localized calculation of SST tendencies allows to decrease the dimensionality of the JPD two times (as the dimension of JPD for both SST and its tendencies is two times larger compared to the JPD only for SST) and thus makes its calculation more efficient without compromising on accuracy. Put differently, sampling SST fields and their associated tendencies jointly would require estimating the full JPD in a space of dimension $2n$, where $n$ is the number of spatial grid points used to discretize the SST field. In contrast, we sample only from the JPD of SST fields (which is of dimension $n$), and then calculate SST tendencies locally. This reduces the required dimensionality of the JPD two times (from $2n$ to $n$), since the tendencies are not independently sampled but are functionally linked to the sampled SST states via local averaging.
@@
It is worth mentioning that tendencies can also be computed in a probabilistic way from a local JPD which is constructed based on the information from the neighborhood of the new tendency (red contour in figure 2), but this method is slower than the one used in this study.
@@
In order to sample from the JPD we use the inverse transform sampling method [10]. We also tried the rejection sampling method but did not gain that much of a difference in its favour; the sampling method based on a reverse-time stochastic differential equation [28] might be another option instead of the inverse transform sampling. Sampling from the JPD allows us to probabilistically reconstruct the originally sparsely-populated reference phase space by filling it with new states and directional vectors (tendencies) needed to advect the image point (figure 1d). Note that the probabilistically-reconstructed reference data can be used not only in the HP method, but in other data-driven methods which suffer a lack of data.
@@
We would like to draw the reader’s attention to the fact that the new states generated by sampling from the JPD does not mean new reference states (i.e. the states generated by the underlying reference model, it is the NEMO model in our case), since sampling from the JPD cannot guarantee that sampled states are reference states, and the HP dynamics in probabilistically-reconstructed phase spaces cannot guarantee it either.
To explain how the sampling procedure calculates a new state, we consider a three-dimensional case consistent with Chua’s circuit – a set of points in an $XYZ$-space. We sample new points from the joint probability distribution $P(x,y,z)$ by decomposing it into a marginal and conditional distributions:
$P(x,y,z)=P_{X}(x)\,P_{Y|X}(y\mid x)\,P_{Z|X,Y}(z\mid x,y).$
First, the marginal distribution $P_{X}(x)$ is estimated from the available data and converted into a cumulative distribution function (CDF), $F_{X}(x)$. A uniform random number $u\in[0,1]$ is then drawn, and the new $x$-coordinate is obtained by inverting the CDF:
$x_{\mathrm{new}}=F_{X}^{-1}(u).$
Next, the conditional distribution $P_{Y|X}(y\mid x_{\mathrm{new}})$ is determined, again converted into a CDF $F_{Y|X}(y)$, and a second uniform random number $u^{\prime}\in[0,1]$ is drawn to produce:
$y_{\mathrm{new}}=F_{Y|X}^{-1}(u^{\prime}).$
Finally, the conditional distribution $P_{Z|X,Y}(z\mid x_{\mathrm{new}},y_{\mathrm{new}})$ is estimated, converted into a CDF $F_{Z|X,Y}(z)$, and a third uniform random number $u^{\prime\prime}\in[0,1]$ is drawn to compute:
$z_{\mathrm{new}}=F_{Z|X,Y}^{-1}(u^{\prime\prime}).$
This sequential inverse transform sampling procedure is repeated to generate as many new points $(x_{\mathrm{new}},y_{\mathrm{new}},z_{\mathrm{new}})$ as required, ensuring that the samples follow the original 3D JPD without explicitly constructing a full 3D JPD, thereby avoiding the higher memory cost associated with direct estimation in three dimensions.
@@
(e) Advection of the image point in probabilistically-reconstructed phase space. In the fifth step, the probabilistically-reconstructed phase space is used by Advection of the image point to calculate the trajectory (red line in figure 1e). For the reader’s convenience, we briefly describe how Advection of the image point works. This method falls into the category of data-driven methods from the hyper-parametersation (HP) class which takes advantage of working in phase spaces, as opposed to the conventional methods operating in the physical space. The method has been tested on both idealized and comprehensive ocean models (two-layer quasi-geostrophic model in a channel and MITgcm in the North Atlantic configuration) and showed significant improvements of the HP solution toward the reference one [19, 22]. The HP approach currently umbrellas five different methods (ranging from purely data-driven to hybrids, which combine the data- and physics-driven paradigms). Another striking feature of the HP approach is that its measure of goodness is how close the HP solution is to the reference phase space. To put it another way, the HP approach matches phase spaces (the reference phase space and the phase space where the HP solution evolves), whereas conventional methods match individual trajectories. This measure allows the HP solution to evolve in the neighborhood of the reference phase space, and therefore reproduce the flow dynamics which is very similar to the reference one.
@@
$\frac{d\mathbf{y}}{dt}=\frac{1}{M}\sum\limits_{j\in\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})+\eta\left(\frac{1}{N}\sum\limits_{i\in\mathcal{U}_{I}}\mathbf{x}_{i}-\mathbf{y}(t)\right),\quad\mathbf{y}(t_{0})=\mathbf{y}_{0},$ (1)
@@
In a nutshell, the actual dynamics is given by the observed reference tendency $\frac{1}{M}\sum\nolimits_{j\in\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})$ the imperfect reconstruction of which from the reference data is compensated by nudging towards the observed reference state $\frac{1}{N}\sum\nolimits_{i\in\mathcal{U}_{I}}\mathbf{x}_{i}$ in the probabilistically-reconstructed phase space. Note that the neighbourhood is computed as the average over $M$ (and $N$ for the nudging term) nearest, in $l_{2}$ norm, to the solution $\mathbf{y}(t)$ points. The choice of the norm and the way the neighborhood $\mathcal{U}(\mathbf{y}(t))$ is calculated is not limited to those used in this study, and can vary depending on what is needed. Our choice is, probably, the simplest, but it serves well the purpose of this work. We dub the set of parameters $\{M,N,\eta\}$ hyper-parameters. The hyper-parameters can be set based on the chosen metric and available data, we will return to their choice later. The interested reader is referred to [19, 22] for a more detailed discussion of the method.
Figure 3: Shown is a sketch of the probabilistically-reconstructed reference phase space (orange blob), vectors $\mathbf{F}(\mathbf{x}_{i})$ (black arrows) pointing out from points $\mathbf{x}_{i}$. For the HP method all these points is a cloud of points which are not ordered in time. The HP solution $\mathbf{y}(t)$ is shown as a red trajectory. The filled red circle denotes the current state $\mathbf{y}(t)$ at time $t$, while the empty red circles represent other points along the HP trajectory. The neighborhood $\mathcal{U}(\mathbf{y}(t))$ is illustrated by the black contour surrounding the filled red circle. The red vectors (and also each point they are attached to) within the neighborhood are the $N$ reference fields $\mathbf{x}_{i},\,i\in\mathcal{U}_{I}$ (and $M$ reference tendencies $\mathbf{F}(\mathbf{x}_{i})$) nearest, in $l_{2}$-norm, to the HP solution $\mathbf{y}(t)$. We assume that $\mathcal{U}_{I}=\mathcal{U}_{J}$ and $M=N$, but the method is not bound to this particular choice.
The characteristic feature of the HP method is that it looks at the reference dataset as being a cloud of points which are not ordered in time in contrast to traditional methods, which rely on the chronological order. It allows the HP method to constantly stay in the reference phase space, as there always exists a neighborhood of nearest points. Thus, the HP method never runs out of reference data. This is why we write $\mathbf{x}_{i}$ and $\mathbf{F}(\mathbf{x}_{j})$ instead of $\mathbf{x}(t)$ and $\mathbf{F}(\mathbf{x}(t))$ in equation (1). At each time step, the HP method takes $N$ nearest $\mathbf{x}_{i}$ (and $M$ tendencies $\mathbf{F}(\mathbf{x}_{j})$) to the solution $\mathbf{y}(t)$ points (figure 3). The main advantage of using the HP method in probabilistically-reconstructed phase spaces is that the data for the method is sampled from the same distribution as the reference data itself.